# CBGBench: Fill in the Blank of Protein-Molecule Complex Binding Graph

### v2

@@ 5.1.2 Result Analysis
LiGAN 0.4645 0.5673 0.0096 0.0718 4.75
3DSBDD 0.5024 0.3904 0.2482 0.8683 6.75
GraphBP 0.5182 0.5645 0.8634 0.9974 9.50
Pocket2Mol 0.5433 0.4922 0.0576 0.4499 6.50
TargetDiff 0.2659 0.3769 0.0483 0.4920 3.00
DiffSBDD 0.3501 0.4588 0.1083 0.6578 5.25
DiffBP 0.3453 0.4621 0.0449 0.4077 3.25
FLAG 0.4215 0.4304 0.6777 0.9769 7.00
D3FG 0.3727 0.4700 0.2115 0.8571 6.50
DecompDiff 0.2576 0.3473 0.0462 0.5248 2.50
@@
LiGAN 1.07 1.20 3.07 1.05 3
3DSBDD 1.07 1.40 1.70 0.65 6
GraphBP 0.13 1.35 0.27 0.10 10
Pocket2Mol 0.67 1.20 2.34 0.70 5
TargetDiff 1.50 1.30 2.77 1.40 1
DiffSBDD 0.70 1.40 1.03 0.95 8
DiffBP 0.23 1.10 1.73 1.35 7
FLAG 0.63 1.25 1.00 0.60 9
D3FG 1.33 1.70 1.33 0.70 4
DecompDiff 1.67 1.40 2.17 1.50 2
Geometry.  From Table. 6, conclusions can be drawn that (i) DecompDiff generates overall the most realistic structures, according to the internal geometries, and TargetDiff is comparable. (ii) The molecule-protein clashes can usually be avoided by diffusion models, like DiffBP, TargetDiff and DecompDiff. However, in the auto-regressive models, only Pocket2Mol avoids clashes. The high frequency of clashes in auto-regressive methods can be explained as follows: These models first identify frontier amino acids or atoms within the molecule. During the placing of new atoms, incorrect binding site localization leads to an erroneous direction for the molecule’s auto-regressively growing path, causing it to extend inward towards the protein. For details, see Appendix. C.1.3.
Conclusion and Discussion.
From Table. 7 and the previous discussion, we conclude that
  • (i)
    The CNN-Based methods like LiGAN are still very competitive, which explains why such methods in the field of drug design still prevail in recent years [50, 51].
  • (ii)
    TargetDiff achieves the best overall performance. In contrast, DecompDiff as the version of TargetDiff that incorporates domain knowledge, D3FG and FLAG show degeneration in performance. This observation reveals that the current incorporation of physicochemical priors can hardly improve the quality of generated molecules. Effectively integrating domain knowledge to guide the model to generate structurally sound and biologically functional molecules remains a challenge. For example, atom clashes are very common in generated molecules, and integrating domain knowledge into general sampling techniques may help mitigate this issue [52, 53].
  • (iii)
    Only Pocket2Mol as an auto-regressive method achieves competitive results, which we attribute to the following reasons: First, it utilizes chemical bonds to constrain atoms to grow orderly along chemical bonds rather than to grow based on distance to the pocket as in DiffBP. Second, it simultaneously predicts the bond types and employs contrastive learning by sampling positive and negative instances of atom positions as real and fake bonds, which is not fully considered by FLAG, enhancing the model’s ability to perceive chemical bond patterns. Therefore, we believe that for autoregressive methods, enabling the model to successfully capture the patterns of chemical bonds is essential.
5.2 Extension on Subtasks
Setup.  Lead optimization is to strengthen the function or property of the binding molecules by remodeling the existing partial 3D graph of molecules. We show the interaction analysis and chemical property in the main context and omit the interaction pattern analysis since maintaining the patterns is not necessary for lead optimization. From the discussion above, the data-driven methods can reach even better performance than the domain-knowledge-based ones, so we have not compared them here. Another reason is that for fragment-based methods, some tasks do not support the direct application of these methods, such as linker design. Besides, the voxelize-based methods are not easily extended to these tasks, and we regard the transferring of methods like LiGAN and 3DSBDD as future work. Hence, we here compare 5 methods that model atoms’ continuous positions with GNNs and have not employed domain knowledge. When training, since the number of training instances is smaller, we use the pretrained models on de novo generation and finetune the auto-regressive models with 1,000,000 iterations. For diffusion-based models with one-shot generation, we train them from scratch because the zero-center-of-mass [54] technique is shifted from employing protein geometric centers to using molecule context’s ones. For detailed results, please refer to Appendix. C.2.
@@ 5.2 Extension on Subtasks
    Overall, the performance gap among these methods in the lead optimization is not as pronounced compared to de novo generation. TargetDiff and Pocket2Mol still perform well. Notably, in the evaluation of initial stability, DiffSBDD, DiffBP, and GraphBP can optimize and complete the binding graph near the given partial molecular conformations, according to the columns of Vina Score and Vina Min. The applicability of GraphBP also reflects the Argument. (ii) in de novo geometry analysis: The failure of GraphBP mainly stems from the difficulty in locating the correct atoms for auto-regressive growth.
@@
    In these tasks, scaffold hopping is the most challenging, as the improvement of generated molecules relative to the reference is minimal; Linker design is relatively the easiest. Notably, in scaffolding, there are still some methods that fail.
Additionally, there are several points that warrant further detailed design. For instance, in linker design, the fragments of the molecule that need to be connected may change orientation due to variations in the linker [18]; In side chain decoration, the protein’s side chains, which are the main components in interaction, should be generated together with the molecule’s side chains to achieve a stable conformation as the entire complex [55, 56].
@@
Linker GraphBP  - 5.63 -0.97 12.43 -7.51 28.18 -7.36 0.3288 0.41 0.86 0.70 3.60 3.92
Pocket2Mol -6.89 18.99 -7.19 25.04 -8.07 37.22 -3.85 0.3276 0.45 1.93 0.67 4.25 2.17
TargetDiff -7.22 36.11 -7.60 41.31 -8.49 50.73 2.61 0.2993 0.39 1.63 0.61 4.02 2.33
DiffSBDD -5.64 11.06 -6.38 19.15 -7.88 34.97 -4.42 0.3110 0.42 1.14 0.66 4.10 3.33
DiffBP -6.27 35.49 -7.19 36.80 -8.74 54.33 6.60 0.3078 0.43 3.45 0.55 4.01 2.42
Fragment GraphBP -5.54 4.86 -6.28 9.78 -7.16 16.21 -11.88 0.3749 0.54 0.87 0.66 4.66 3.00
Pocket2Mol -6.87 22.78 -7.61 37.45 -8.33 54.05 -0.28 0.3310 0.46 1.02 0.63 4.07 1.75
TargetDiff -6.06 24.56 -6.78 30.43 -7.96 42.00 -2.38 0.3003 0.45 1.43 0.58 4.28 2.67
DiffSBDD -4.64 19.14 -5.84 28.90 -7.66 37.18 -6.67 0.3076 0.47 0.73 0.58 4.39 3.42
DiffBP -4.51 22.31 -6.18 29.52 -7.90 45.70 -1.92 0.2952 0.46 2.24 0.49 4.30 3.33
Side chain GraphBP -5.01 10.15 -5.43 11.46 -6.14 9.71 -11.05 0.4459 0.61 1.93 0.76 4.93 3.17
Pocket2Mol -5.99 22.26 -6.56 33.29 -7.26 41.04 -4.34 0.3600 0.49 0.21 0.65 4.20 2.25
TargetDiff -5.80 23.90 -6.50 35.81 -7.40 46.87 -2.55 0.3213 0.48 0.88 0.60 4.41 2.08
DiffSBDD -4.43 15.12 -5.99 30.23 -7.58 44.09 -9.38 0.3178 0.43 1.20 0.65 4.03 3.17
DiffBP -4.61 14.31 -5.73 24.29 -7.03 38.96 -7.38 0.3143 0.49 1.29 0.56 4.50 3.50
Scaffold GraphBP  - 0.00  - 0.06 -3.90 0.99 -50.62 0.3797 0.43 0.14 0.76 4.98 3.58
Pocket2Mol -4.80 16.84 -5.71 23.08 -6.89 38.18 -8.07 0.3378 0.43 0.91 0.64 4.48 2.00
TargetDiff -5.52 31.47 -5.86 34.39 -7.06 44.32 -6.22 0.3038 0.43 0.89 0.59 4.26 1.83
DiffSBDD -3.85 18.44 -4.90 22.12 -6.81 34.98 -10.23 0.2985 0.42 -0.13 0.53 4.29 3.33
DiffBP -2.09 13.89 -4.35 16.84 -6.46 32.43 -12.14 0.3025 0.43 3.37 0.56 4.44 3.42
5.3 Case study
@@
Setup.  On these targets, there are molecules reported active to them experimentally, we here randomly select 200 of them for each target and conduct two kinds of experiments. Firstly, we try to figure out if the model can generate binding molecules that have similar chemical distributions with the actives. We use extended connectivity fingerprint (ECFP[57] to get the molecule fingerprint and employ t-SNE [58] for 2-dimensional visualization. Secondly, we aim to find out the distribution of Vina Docking Energy and LBE of the generated and the actives as the metrics for binding affinities. Seven methods except for voxelized-based ones (inflexible to extend to the framework) and DecompDiff (requiring complex data-preprocessing for domain knowledge and differing from TargetDiff slightly) are tested on the two targets. Besides, we select 200 molecules randomly from GEOM-DRUG [59], as a randomized control sample set.
@@
    In Figure. 3, we can see that Pocket2Mol, TargetDiff, and D3FG have better consistency in the chemical distribution of molecules, as evidenced by a greater degree of overlap with the actives. In addition, in comparison to randomly selected molecules in GEOM-DRUG, these models show different preferences in generating binding molecules since the clustering center in the chemical space differs.
@@
    Figure. 4 shows that in generating molecules based on the real-world target ADRB1, D3FG exhibits superior and stable performance. TargetDiff also perform well.
These conclusions are essentially consistent with the conclusion in Sec. 5.1, reflecting that the established evaluation protocols exhibit strong consistency and generalizability on real-world target data. Besides, it is worth noting that DiffBP, FLAG, and Pocket2Mol can possibly generate molecules with small atom numbers and high LBE. This indicates that they have the potential to excel in lead discovery on ADRB1, as a good lead should possess good synthesizability and modifiability, smaller molecular weight, and stable binding conformations. For DRD3, please refer to Appendix. C.2.1 for details.
@@ 6 Conclusion and Limitation
Appendix A Supplementary Evaluation Details
A.1 Included Functional Groups

### v3

@@ 5.2 Extension on Subtasks
LiGAN 0.4645 0.5673 0.0096 0.0718 5.75
3DSBDD 0.5024 0.3904 0.2482 0.8683 8.75
GraphBP 0.5182 0.5645 0.8634 0.9974 11.50
Pocket2Mol 0.5433 0.4922 0.0576 0.4499 8.50
TargetDiff 0.2659 0.3769 0.0483 0.4920 4.50
DiffSBDD 0.3501 0.4588 0.1083 0.6578 7.25
DiffBP 0.3453 0.4621 0.0449 0.4077 5.25
FLAG 0.4215 0.4304 0.6777 0.9769 9.00
D3FG 0.3727 0.4700 0.2115 0.8571 8.50
DecompDiff 0.2576 0.3473 0.0462 0.5248 4.00
MolCraft 0.2250 0.2683 0.0264 0.2691 2.00
VoxBind 0.2701 0.3771 0.0103 0.1890 3.00
@@
LiGAN 1.13 1.45 3.63 1.25 3
3DSBDD 1.13 1.70 1.93 0.65 7
GraphBP 0.17 1.65 0.27 0.10 12
Pocket2Mol 0.73 1.45 2.53 0.70 7
TargetDiff 1.73 1.55 2.93 1.50 2
DiffSBDD 0.77 1.70 1.10 0.95 10
DiffBP 0.27 1.35 1.83 1.35 9
FLAG 0.70 1.55 1.20 0.60 11
D3FG 1.47 2.00 1.53 0.70 6
DecompDiff 1.90 1.65 2.57 1.60 4
MolCraft 1.80 1.60 3.30 2.00 1
VoxBind 1.40 1.85 3.20 1.80 5
@@
    Overall, the performance gap among these methods in the lead optimization is not as pronounced compared to de novo generation. MolCraft, Target and Pocket2Mol maintain the good performance. Notably, in the evaluation of the stability of initial poses, the other three also complete the binding graph near the given partial molecular conformations well, according to the columns of Vina Score and Vina Min. The applicability of GraphBP also reflects the Argument. (ii) in our de novo geometry analysis: The failure of GraphBP mainly stems from the difficulty in locating the correct atoms for auto-regressive growth.
@@
    In these tasks, scaffold hopping is the most challenging task, as the improvement of generated molecules relative to the reference is minimal; Linker design is relatively the easiest. Notably, in scaffolding, there are still some methods that fail.
  • (iii)
    There is a large space for improvements in these tasks, since the $\mathrm{MPBG}\%$ metrics are usually negative, indicating that in most cases, the molecule is not optimized. Some edge-cutting techniques such as DPO (Rafailov et al., 2024; Cheng et al., 2024; Gu et al., 2024) and ITA (Yang et al., 2020; Kong et al., 2023; Lin et al., 2024) may be used to augment the optimized molecules as supervision signals for the models.
Additionally, several points warrant further detailed design. For instance, in linker design, the fragments of the molecule that need to be connected may change orientation due to variations in the linker (Guan et al., 2023a); In side chain decoration, the protein’s side chains, which are the main components in interaction, should be generated together with the molecule’s side chains to achieve a stable conformation as the entire complex (Luo et al., 2023; Huang et al., 2024b).
5.3 Case study on Real-World Targets
Setup.  On these targets, there are molecules reported active to them experimentally, we here randomly select 200 of them for each target and conduct two kinds of experiments. Firstly, we try to figure out if the model can generate binding molecules that have similar chemical distributions with the actives. We use extended connectivity fingerprint (ECFP(Rogers & Hahn, 2010) to get the molecule fingerprint and employ t-SNE (van der Maaten & Hinton, 2008) for 2-dimensional visualization. Secondly, we aim to find out the distribution of Vina Docking Energy and LBE of the generated and the actives as the metrics for binding affinities. Eight methods except for voxelized-based ones (inflexible to extend) and DecompDiff (requiring complex data-preprocessing for domain knowledge) are tested on the two targets. Besides, we select 100 molecules randomly from GEOM-DRUG (Axelrod & Gómez-Bombarelli, 2022), as a randomized control sample set.
@@ 5.3 Case study on Real-World Targets
Linker GraphBP  - 5.63 -0.97 12.43 -7.51 28.18 -7.36 0.3288 0.41 0.86 0.70 3.60 4.75
Pocket2Mol -6.89 18.99 -7.19 25.04 -8.07 37.22 -3.85 0.3276 0.45 1.93 0.67 4.25 2.83
TargetDiff -7.22 36.11 -7.60 41.31 -8.49 50.73 2.61 0.2993 0.39 1.63 0.61 4.02 3.17
DiffSBDD -5.64 11.06 -6.38 19.15 -7.88 34.97 -4.42 0.3110 0.42 1.14 0.66 4.10 4.25
DiffBP -6.27 35.49 -7.19 36.80 -8.74 54.33 6.60 0.3078 0.43 3.45 0.55 4.01 3.25
MolCraft -7.13 38.80 -7.81 42.56 -8.81 58.44 5.09 0.3334 0.43 1.06 0.67 4.14 1.50
Fragment GraphBP -5.54 4.86 -6.28 9.78 -7.16 16.21 -11.88 0.3749 0.54 0.87 0.66 4.66 3.58
Pocket2Mol -6.87 22.78 -7.61 37.45 -8.33 54.05 -0.28 0.3310 0.46 1.02 0.63 4.07 2.00
TargetDiff -6.06 24.56 -6.78 30.43 -7.96 42.00 -2.38 0.3003 0.45 1.43 0.58 4.28 3.33
DiffSBDD -4.64 19.14 -5.84 28.90 -7.66 37.18 -6.67 0.3076 0.47 0.73 0.58 4.39 4.17
DiffBP -4.51 22.31 -6.18 29.52 -7.90 45.70 -1.92 0.2952 0.46 2.24 0.49 4.30 4.08
MolCraft -6.75 21.12 -7.06 36.07 -7.92 43.02 -0.09 0.3236 0.46 1.27 0.51 4.64 2.58
Side chain GraphBP 5.01 10.15 -5.43 11.46 -6.14 9.71 -11.05 0.4459 0.61 1.93 0.76 4.93 3.75
Pocket2Mol -5.99 22.26 -6.56 33.29 -7.26 41.04 -4.34 0.3600 0.49 0.21 0.65 4.20 2.91
TargetDiff -5.80 23.90 -6.50 35.81 -7.40 46.87 -2.55 0.3213 0.48 0.88 0.60 4.41 2.58
DiffSBDD -4.43 15.12 -5.99 30.23 -7.58 44.09 -9.38 0.3178 0.43 1.20 0.65 4.03 3.83
DiffBP -4.61 14.31 -5.73 24.29 -7.03 38.96 -7.38 0.3143 0.49 1.29 0.56 4.50 4.25
MolCraft -6.10 24.10 -6.64 35.58 -7.49 41.67 -3.12 0.3227 0.44 1.22 0.61 4.36 2.42
Scaffold GraphBP  - 0.00  - 0.06 -3.90 0.99 -50.62 0.3797 0.43 0.14 0.76 4.98 4.16
Pocket2Mol -4.80 16.84 -5.71 23.08 -6.89 38.18 -8.07 0.3378 0.43 0.91 0.64 4.48 2.42
TargetDiff -5.52 31.47 -5.86 34.39 -7.06 44.32 -6.22 0.3038 0.43 0.89 0.59 4.26 2.00
DiffSBDD -3.85 18.44 -4.90 22.12 -6.81 34.98 -10.23 0.2985 0.42 -0.13 0.53 4.29 4.17
DiffBP -2.09 13.89 -4.35 16.84 -6.46 32.43 -12.14 0.3025 0.43 3.37 0.56 4.44 4.17
MolCraft -4.71 14.03 -5.35 30.16 -7.02 43.53 -7.34 0.3146 0.42 1.01 0.56 4.55 2.83
@@
    In Figure. 4, we can see that Pocket2Mol, TargetDiff, D3FG, and MolCraft have better consistency in the chemical distribution of molecules, as evidenced by a greater degree of overlap with the actives. In addition, in comparison to randomly selected molecules in GEOM-DRUG, these models show different preferences in generating binding molecules since the clustering center in the chemical space differs.
@@
    Figure. 5 shows that in generating molecules based on the real-world target ADRB1, D3FG exhibits superior performance. TargetDiff and MolCraft perform comparably.
These conclusions are essentially consistent with the conclusion in Sec. 5.1, reflecting that the established evaluation protocols exhibit strong consistency and generalizability on real-world target data. Besides, it is worth noting that DiffBP, GraphBP, and Pocket2Mol can possibly generate molecules with small atom numbers and high LBE. This indicates that they have the potential to excel in lead discovery on ADRB1, as a good lead should possess good synthesizability and modifiability, smaller molecular weight, and stable binding conformations. For DRD3, please refer to Appendix. E.3 for details.
@@ 6 Conclusion and Limitation
Appendix A Review on the Methods’ Timeline
Here we give a brief review of the 12 included methods in chronological order, with Table. LABEL:tab:appreview giving some standardized information on these methods.
  • LiGan is firstly proposed to incorporate a generative model for the SBDD task, in which the encoder and the decoder are 3-layer CNNs, and the Variational Auto-Encoder is employed to model the probability of the atom’s density map in a molecule. It split the training and test set in CrossDocked2020.
  • Due to the rise of Graph Neural Networks (GNNs), 3DSBDD subsequently adopted GNNs for modeling this task. At that time, GNNs were more adept at handling pairwise information, so an autoregressive graph network modeling approach was chosen as an effective method for molecular modeling. However, the development of graph networks for 3D tasks remained slow, leading 3DSBDD to continue modeling molecules using voxelized maps. 3DSBDD leverages graph networks to predict the voxel position that the next atom will occupy, as well as the atom type, thereby generating molecules in an autoregressive manner. In addition, 3DSBDD further standardized the evaluation protocol of the experiment by using Vina Docking Energy to assess the stability of the generated molecules.
  • The rise of EGNN has enabled autoregressive modeling of continuous atomic coordinates in the SBDD task while preserving equivariance and invariance properties. Following this development, GraphBP adopted the EGNN architecture, using a depth-first search approach to rank the molecules from near to far on the protein surface. This allows for autoregressive molecular generation starting from the protein frontier and ultimately modeling a complete molecule. However, GraphBP did not consider critical chemical priors such as bond lengths and bond angles between atoms, making it challenging to capture the internal patterns of molecules. The evaluation is based on Gnina Docking, different from 3DSBDD.
  • Meanwhile, Pocket2Mol directly modeled atomic coordinates using GVP. Compared to GraphBP, Pocket2Mol took molecular bonds into account and employed contrastive learning, where false atoms were used to regularize the coordinates and bonding of real atoms. This significantly improved the stability of the generated molecules. The evaluation follows 3DSBDD.
  • The rise of generative models, particularly the widespread application of diffusion models in the field of images, has drawn attention to these cutting-edge generative techniques in the molecular domain as well. Diffusion models are inspired by particle systems in statistical physics, which aligns closely with the task of generating atomic coordinates in molecules. DiffSBDD, as pioneering works in applying diffusion models to the SBDD task, have also achieved significant results. DiffSBDD followed most of the concepts from EDM, such as the zero-center-of-mass technique, and was tested and generalized on Binding MOAD beyond CrossDocked2020.
  • Meanwhile, DiffBP also proposed to use the Gaussian Diffusion model on atoms’ positions and masked type diffusion for elements’ types. DiffBP analyzed the drawbacks of autoregressive models from the perspective of the Boltzmann energy distribution in statistical physics and observed that molecular weight severely affects docking performance, being the first to adopt a grouping approach for model comparison.
  • After that, following previous explorations like GeoDiff and EDM in molecular (conformation) generation, TargetDiff innovatively introduced a diffusion model-based SBDD method, achieving new state-of-the-art (SOTA) performance. For validation, TargetDiff first proposed using Vina Score and Vina Min mode to assess the stability of the generated molecules’ initial states. Additionally, the embeddings learned by the model can also be used for affinity prediction.
  • On the other hand, fragment-based generative methods are also in the early stages of exploration. FLAG follows the autoregressive generation approach of Pocket2Mol, breaking molecules into fragments and individual atoms, allowing for the generation of complex molecules with fewer steps. However, its modeling approach lacks sufficient consideration of chemical bond participation, which is a key reason for its suboptimal molecular generation performance.
  • Afterward, D3FG combined fragment-based generation with the diffusion method, leveraging protein generation techniques to establish a diffusion model that predicts the orientation matrix, central coordinates, and fragment types based on fragment information. The model then uses linkers to connect the molecule fragments. Additionally, D3FG enables molecular optimization by replacing fragments that are either prominent or insufficiently contributing.
  • In contrast to D3FG, which treats fragments as rigid bodies, DecompDiff considers fragments as flexible structures, thereby decomposing molecules into arms and scaffolds, and performing multivariate Gaussian diffusion within different decomposition clusters. Additionally, DecompDiff employs non-intersection guidance to mitigate the issue of atomic overlap between the molecule and protein amino acids.
  • In addition, work on voxelized maps is steadily progressing. VoxBind builds on the pioneering work of VoxMol, using diffusion-based denoising modeling on the voxelized density map of molecules. During the generation steps, it incorporates wj-sampling to reduce computational complexity, while also producing structurally stable molecules.
  • More recently, MolCraft was proposed, utilizing the Bayes Flow Network, which is designed for modeling probabilistic densities’ parameters. This allows molecules to avoid the difficulty of coupling the two modalities: the discrete variable of atomic types and the continuous positions of atoms in the molecule. It effectively addresses issues such as mode collapse and has demonstrated outstanding performance as the most recent SOTA model.
Table 9: A brief review of the included methods.
Method Time Generative Model Network Structure Prior Knowledge Evaluation Datasets
LiGan Oct. 2020 VAE CNN None CrossDocked2020
3DSBDD Mar. 2022 Auto-regressive GNN None CrossDocked2020
GraphBP Apr. 2022 Auto-regressive EGNN None CrossDocked2020
Pocket2Mol May. 2022 Auto-regressive GVP None CrossDocked2020
DiffSBDD Oct. 2022 DDPM EGNN None
CrossDocked2020
Binding MOAD
DiffBP Nov. 2022 DDPM GVP None CrossDocked2020
TargetDiff Feb. 2023 DDPM EGNN None CrossDocked2020
FLAG Feb. 2023 Auto-regressive GVP Fragment CrossDocked2020
D3FG May. 2023 DDPM EGNN + IPA Fragment CrossDocked2020
DecompDiff Feb. 2024 DDPM EGNN Arm&Scaffold CrossDocked2020
MolCraft Apr. 2024 BFN EGNN None CrossDocked2020
VoxBind May. 2024 DDPM CNN None CrossDocked2020
Appendix B Supplementary Evaluation Details
B.1 Included Functional Groups